A student is chosen a random from a class of 20 boys and 15 girls. Find the odds a girl is chosen.
Answer by stanbon(75887)

A student is chosen a random from a class of 20 boys and 15 girls. Find the odds a girl is chosen.
---------------------------------
P(girl) = 15/35
P(boy) = 20/35
---
Odds in favor of girl = P(girl)/P(boy) = 15:20
================================================
Cheers,
Stan H.
===============